### 07 Jan 2024 v5

@@ IV-B Benchmarks and Metrics
To thoroughly assess the generalization capability of the learned image compression models, we conduct performance evaluations on three distinct datasets, including Kodak [55], Tecnick [56], CLIC Professional Valid [22].
@@
  • Kodak [55] is selected as a test set for almost all end-to-end image compression models [15, 16, 17, 19, 24, 23, 31, 14, 6, 28, 32, 33, 25, 29, 34, 30, 35, 27]. It contains 24 raw $768\times 512$ images.
@@
  • Tecnick [56] contains 100 $1200\times 1200$ RGB images, which is used as a test set in many methods [17, 19, 28, 49, 30, 36, 27, 38, 35, 59].
@@
  • CLIC Professional Valid [22] is the validation set of 3rd Challenge on Learned Image Compression which contains $41$ images. Image in this dataset contains around $2048\times 1440$ pixels. CLIC Pro Valid is widely used in many recent methods [24, 28, 29, 25, 27, 30, 57, 58, 35].
@@
We use the Bjøntegaard delta rate (BD-Rate) [68] to evaluate the performance of learned image compression models.
IV-C Rate-Distortion Performance
IV-C1 Quantitive Results
@@
Rate-distortion curves are presented in Fig. 10. When compared with Cheng’20 [24], our proposed ${}^{++}$ can achieve a maximum improvement of $0.5\sim 0.8$ dB in PSNR and achieve a maximum improvement of $0.6$ dB in MS-SSIM in dB. Our ${}^{++}$ adopts simplified analysis transform and synthesis transform of Cheng’20 [24], therefore, the improvement of model performance is attributed to our linear complexity multi-reference entropy modeling. Our linear complexity attention-based multi-reference entropy models can capture more contexts, which leads to much better rate-distortion performance. The improvement also proves correlations exist in multiple dimensions since Cheng’20 [24] adopts an spatial auto-regressive context module. Compared with ELIC [29], our ${}^{++}$ can be up to $0.4$ dB higher at low bit rates in PSNR and $1$ dB higher in MS-SSIM [61].
@@
BD-rate reductions are presented in Table III. When computing BD-rate, VTM-17.0 Intra under YUV444 is employed as anchor. Unofficial weights of ELIC [^2] https://github.com/VincentChandelier/ELiC-ReImplemetation are used to evaluate the rate-distortion performance of ELIC on Tecnick, CLIC Professional Valid. Our ${}^{++}$ outperforms our previous MLIC and ${}^{+}$ [30]. Our ${}^{++}$ reduces BD-rate by $13.39\%$ over VTM-17.0 Intra on Kodak while existing method ELIC only reduces $5.95\%$ BD-rate and LIC-TCM only reduces $10.14\%$ BD-rate on Kodak. Moreover, our ${}^{++}$ performs better on high-resolution datasets, such as Tecnick and CLIC Professional Valid. Our ${}^{++}$ reduces $17.59\%$ on Tecnick and reduces $13.08\%$ BD-rate on CLIC Professional Valid, which are much better than existing methods. Our ${}^{++}$ achieves state-of-the-art performance on all three datasets. The excellent performance of our ${}^{++}$ on these datasets also demonstrates the excellent generalization of our ${}^{++}$. We highlight the BD-rate for MS-SSIM, our proposed ${}^{++}$ reduces about $50\%$ bits compared to VTM-17.0 Intra, which is a large progress in learned image compression.
IV-C2 Qualitive Results
@@
To further demonstrate superiority of our proposed ${}^{++}$, we compare our ${}^{++}$ with learned image compression models Cheng’20 [24], Xie’21 [28], STF [25], WACNN [25], ELIC [29] and non-neural codec VTM-17.0 Intra [5] on perceptual quality. Fig. 11 presents the reconstructions of Kodim07 from Kodak. PSNR value of the image reconstructed by our ${}^{++}$ optimized for MSE is $1$dB higher than image reconstructed by VTM-17.0 Intra [5]. MS-SSIM of the image reconstructed by our ${}^{++}$ optimized for MS-SSIM is $0.02$ higher than image reconstructed by VTM-17.0 Intra. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of ${}^{++}$ have sharper textures retain more details. In terms of visual quality, our ${}^{++}$ have significant improvements on rate-perception performance compared to other models.
@@
Figure 12: GPU memory consumption, encoding time, decoding time comparisons among our proposed models [24, 28, 41, 25, 29, 27] and recent learned image compression models. Encoding and decoding time of Qian’21 [26], Entroformer [41], NeuralSyntax [57] exceed 4.5s.
@@
TABLE IV: Ablation Studies on Kodak. BD-rate (%) is employed to evaluate their contributions. ${}^{++}$ is the anchor.
IV-D Computational Complexity
@@
Figure 13: Intra-slice global spatial attention map $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i}_{na,q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i}_{ac,k}\right)^{\top}$ of Kodim01, Kodim11, Kodim20, Kodim21 from Kodak [55] dataset, where $i$ is the index of selected slice, $j$ is the index of selected query. Because interactions within anchor and non-anchor part are not allowed, the attention map is checkerboard-like.
In order to better fit the real-world scenario, we test the model complexities in the case of different resolution images as inputs. We select $16$ images with resolution larger than $3584\times 3584$ from LIU4K test dataset [69] and we center crop these images to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072,3584\times 3584\}$ patches. We compared our ${}^{++}$ with our prior work ${}^{+}$ which employs vanilla attention to illustrate the advantages of proposed linear complexity global context capturing. We also compare our proposed ${}^{++}$ with recent learned image compression models [24, 26, 28, 41, 57, 25, 29, 27]. The experiments are conducted on a Tesla A100 GPU and a Xeon(R) Platinum 8260C CPU. The results are illustrated in Fig. 12.
IV-D1 On GPU Memory comsuption
The quadratic complexity of vanilla attention leads to significantly more memory consumptions on high resolution image coding. When compressing $2048\times 2048$ images, ${}^{+}$ consumes nearly $22.37$ GB GPU memory. Our proposed linear complexity global context capturing significantly reduces the consumption of GPU memory as our proposed ${}^{++}$ only takes about $5.55$ GB GPU memory to compress a $2048\times 2048$ image. When compressing $3072\times 3072$ images, ${}^{+}$ consumes $46.38$ GB GPU memory while our proposed ${}^{++}$ only consumes $7.69$ GB GPU memory. Compared with recent LIC-TCM [27], our ${}^{++}$ consumes $\frac{1}{2}$ of the GPU memory consumed by LIC-TCM when compressing $2048\times 2048$ images. When compressing $2560\times 2560$ images, our ${}^{++}$ only consumes about $\frac{1}{4}$ of the GPU memory consumed by LIC-TCM. When compressing a $3584\times 3584$ image, the GPU memory consumption of our proposed ${}^{++}$ is still only $12.3$ GB while the LIC-TCM requires $45.95$ GB GPU memory. The curve of the GPU memory consumed by our ${}^{++}$ as the resolution grows is also much flatter.
IV-D2 On Encoding and Decoding Time
@@
IV-D3 On Forward MACs
Compared with ${}^{+}$, our ${}^{++}$, which employs the proposed linear complexity global spatial context modules, has lower MACs. Compared with the recent LIC-TCM [27], our ${}^{++}$ demonstrates significantly reduced MACs. Compared with ELIC [29], STF [25], and WACNN [25], our ${}^{++}$ has higher MACs when the input is high-resolution image. One contributing factor to the higher MACs is the utilization of Cheng’20 [24] as the basis for the transform module in our proposed ${}^{++}$. While the context module in Cheng’20 is simpler than that of ELIC, STF, and WACNN, the overall MACs of Cheng’20 are higher due to its transform modules having higher MACs. Since our ${}^{++}$ employs a modification of analysis transform and synthesis transform of Cheng’20, it results in higher MACs compared to ELIC and WACNN. In addition, since our ${}^{++}$ employed an more advanced entropy model, which leads to higher MACs. However, the entropy model has modest effect on the overall MACs since the input image is down-sampled for four times in analysis transform, which implies that designing more advanced entropy models is more resource-efficient.
IV-E Ablation Studies
IV-E1 Settings
Figure 14: Inter-slice global spatial attention map $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i}_{q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i}_{k}\right)^{\top}$ of Kodim01, Kodim11, Kodim20, Kodim21 from Kodak [55] dataset, where $i$ is the index of selected slice, $j$ is the index of selected query.
We conduct corresponding ablation studies and evaluate the contribution of each module on Kodak [55] dataset. Each model is optimized for MSE. We train each model for 2M steps. We use the training strategy in Section IV-A.
IV-E2 Analysis of Channel-wise Context Module
@@
IV-E3 Analysis of Local Spatial Context Module
@@
The vanilla checkerboard context module leads to slight performance degradation [40] compared with pixel-cnn-like serial context modules [39, 19]. The vanilla checkerboard context module contains one convolutional layer which is linear. The other drawback is the fixed kernel weights during inference. In contrast, our proposed checkerboard attention-based local spatial context module is non-linear and incorporates dynamic attention map generation with two-pass decoding, allowing for improved flexibility and adaptability. Our proposed checkerboard attention-based local spatial context module achieves a further reduction of $2.32\%$ in bit-rate compared to vanilla checkerboard context module. Compared with Case 7, which solely incorporates channel-wise context modules, models incorporating both local spatial and channel context modules demonstrate superior performance, further validating the presence of redundancy in the local spatial domain.
IV-E4 Analysis of Intra-Slice Global Context Module
@@
The $i$-th slice is used as an example to illustrate the process. In practice, the computation of $\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}\hat{\bm{y}}^{i}_{ac,v}$ in Equation 15 is performed first for linear complexity. However, we can still compute $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i}_{na,q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i}_{ac,k}\right)^{\top}$ as the attention map to validate the ability to capture global dependencies since the $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i}_{na,q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i}_{ac,k}\right)^{\top}$ is employed as the implicit similarity metric, where $j$ is the index of selected query. The attention maps of Kodim01, Kodim11, Kodim 20, Kodim21 captured by proposed intra-slice global spatial context module $g_{gc,intra}$ is illustrated in Fig. 13. The checkerboard-like pattern in the attention map arises due to the absence of interactions within the anchor and non-anchor parts. Our model successfully captures distant correlations between the anchor and non-anchor parts, which local context modules are unable to achieve. Although our intra-slice global context module may bear some resemblance to cross-attention models, we focus solely on the interactions within a single slice. We only use the attention map of $\hat{\bm{y}}^{i-1}$ to predict correlations in $\hat{\bm{y}}^{i}$. When our proposed global context modules collaborate with local spatial context modules, the overall performance is further improved, underscoring the necessity of global spatial context modules for capturing global correlations and local spatial context modules for capturing local correlations.
IV-E5 Analysis of Inter-Slice Global Context Module

### 20 Feb 2024 v9

@@ 4.2 Benchmarks and Metrics
To thoroughly assess the generalization capability of the learned image compression models, we conduct performance evaluations on three distinct datasets, including Kodak (Kodak, 1993), Tecnick (Asuni & Giachetti, 2014), CLIC Professional Valid (Toderici et al., 2020).
@@
  • Kodak (Kodak, 1993) is selected as a test set for almost all end-to-end image compression models (Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Cheng et al., 2020; Minnen & Singh, 2020; Chen et al., 2021; Guo et al., 2022; Wu et al., 2022; Xie et al., 2021; Gao et al., 2021; Chen et al., 2022; Zou et al., 2022; He et al., 2022; Koyuncu et al., 2022; Jiang et al., 2023b; Duan et al., 2023; Liu et al., 2023). It contains 24 raw $768\times 512$ images.
@@
  • Tecnick (Asuni & Giachetti, 2014) contains 100 $1200\times 1200$ RGB images, which is used as a test set in many methods (Ballé et al., 2018; Minnen et al., 2018; Xie et al., 2021; Kim et al., 2022; Jiang et al., 2023b, a; Liu et al., 2023; Zhu et al., 2022b; Duan et al., 2023; Pan et al., 2022).
@@
  • CLIC Professional Valid (Toderici et al., 2020) is the validation set of 3rd Challenge on Learned Image Compression which contains $41$ images. Image in this dataset contains around $2048\times 1440$ pixels. CLIC Pro Valid is widely used in many recent methods (Cheng et al., 2020; Xie et al., 2021; He et al., 2022; Zou et al., 2022; Liu et al., 2023; Jiang et al., 2023b; Wang et al., 2022; Zhu et al., 2022a; Duan et al., 2023).
@@
We use the Bjøntegaard delta rate (BD-Rate) (Bjontegaard, 2001) to evaluate the performance of learned image compression models.
4.3 Rate-Distortion Performance
4.3.1 Quantitive Results
@@
Rate-distortion curves are presented in Fig. 10. When compared with Cheng’20 (Cheng et al., 2020), our proposed ${}^{++}$ can achieve a maximum improvement of $0.5\sim 0.8$ dB in PSNR and achieve a maximum improvement of $0.6$ dB in MS-SSIM in dB. Our ${}^{++}$ adopts simplified analysis transform and synthesis transform of Cheng’20 (Cheng et al., 2020), therefore, the improvement of model performance is attributed to our linear complexity multi-reference entropy modeling. Our linear complexity multi-reference entropy models can capture more contexts, which leads to much better rate-distortion performance. The improvement also proves correlations exist in multiple dimensions since Cheng’20 (Cheng et al., 2020) adopts an spatial auto-regressive context module. Compared with ELIC (He et al., 2022), our ${}^{++}$ can be up to $0.4$ dB higher at low bit rates in PSNR and $1$ dB higher in MS-SSIM (Wang et al., 2003).
@@
BD-rate reductions are presented in Table 3. When computing BD-rate, VTM-17.0 Intra under YUV444 is employed as anchor. Unofficial weights of ELIC [^2] https://github.com/VincentChandelier/ELiC-ReImplemetation are used to evaluate the rate-distortion performance of ELIC on Tecnick, CLIC Professional Valid. Our ${}^{++}$ outperforms our previous MLIC and ${}^{+}$ (Jiang et al., 2023b). Our ${}^{++}$ reduces BD-rate by $13.39\%$ over VTM-17.0 Intra on Kodak while existing method ELIC only reduces $5.95\%$ BD-rate and LIC-TCM only reduces $10.14\%$ BD-rate on Kodak. Moreover, our ${}^{++}$ performs better on high-resolution datasets, such as Tecnick and CLIC Professional Valid. Our ${}^{++}$ reduces $17.59\%$ on Tecnick and reduces $13.08\%$ BD-rate on CLIC Professional Valid, which are much better than existing methods. Our ${}^{++}$ achieves state-of-the-art performance on all three datasets. The excellent performance of our ${}^{++}$ on these datasets also demonstrates the excellent generalization of our ${}^{++}$. We highlight the BD-rate for MS-SSIM, our proposed ${}^{++}$ reduces about $50\%$ bits compared to VTM-17.0 Intra, which is a large progress in learned image compression.
4.3.2 Qualitive Results
@@
To further demonstrate superiority of our proposed ${}^{++}$, we compare our ${}^{++}$ with learned image compression models Cheng’20 (Cheng et al., 2020), Xie’21 (Xie et al., 2021), STF (Zou et al., 2022), WACNN (Zou et al., 2022), ELIC (He et al., 2022) and non-neural codec VTM-17.0 Intra (Bross et al., 2021) on perceptual quality. Fig. 11 presents the reconstructions of Kodim07 from Kodak. PSNR value of the image reconstructed by our ${}^{++}$ optimized for MSE is $1$dB higher than image reconstructed by VTM-17.0 Intra (Bross et al., 2021). MS-SSIM of the image reconstructed by our ${}^{++}$ optimized for MS-SSIM is $0.02$ higher than image reconstructed by VTM-17.0 Intra. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of ${}^{++}$ have sharper textures retain more details. In terms of visual quality, our ${}^{++}$ have significant improvements on rate-perception performance compared to other models.
@@
Figure 12: GPU memory consumption, encoding time, decoding time comparisons among our proposed models (Cheng et al., 2020; Xie et al., 2021; Qian et al., 2022; Zou et al., 2022; He et al., 2022; Liu et al., 2023) and recent learned image compression models. Encoding and decoding time of Qian’21 (Qian et al., 2020), Entroformer (Qian et al., 2022), NeuralSyntax (Wang et al., 2022) exceed 4.5s.
@@
Table 4: Ablation Studies on Kodak. BD-rate (%) is employed to evaluate their contributions. ${}^{++}$ is the anchor.
4.4 Computational Complexity
@@
Figure 13: Intra-slice global spatial attention map $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i}_{na,q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i}_{ac,k}\right)^{\top}$ of Kodim01, Kodim11, Kodim20, Kodim21 from Kodak (Kodak, 1993) dataset, where $i$ is the index of selected slice, $j$ is the index of selected query. Because interactions within anchor and non-anchor part are not allowed, the attention map is checkerboard-like.
In order to better fit the real-world scenario, we test the model complexities in the case of different resolution images as inputs. We select $16$ images with resolution larger than $3584\times 3584$ from LIU4K test dataset (Liu et al., 2020a) and we center crop these images to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072,3584\times 3584\}$ patches. We compared our ${}^{++}$ with our prior work ${}^{+}$ which employs vanilla attention to illustrate the advantages of proposed linear complexity global context capturing. We also compare our proposed ${}^{++}$ with recent learned image compression models (Cheng et al., 2020; Qian et al., 2020; Xie et al., 2021; Qian et al., 2022; Wang et al., 2022; Zou et al., 2022; He et al., 2022; Liu et al., 2023). The experiments are conducted on a Tesla A100 GPU and a Xeon(R) Platinum 8260C CPU. The results are illustrated in Fig. 12.
4.4.1 On GPU Memory comsuption
The quadratic complexity of vanilla attention leads to significantly more memory consumptions on high resolution image coding. When compressing $2048\times 2048$ images, ${}^{+}$ consumes nearly $22.37$ GB GPU memory. Our proposed linear complexity global context capturing significantly reduces the consumption of GPU memory as our proposed ${}^{++}$ only takes about $5.55$ GB GPU memory to compress a $2048\times 2048$ image. When compressing $3072\times 3072$ images, ${}^{+}$ consumes $46.38$ GB GPU memory while our proposed ${}^{++}$ only consumes $7.69$ GB GPU memory. Compared with recent LIC-TCM (Liu et al., 2023), our ${}^{++}$ consumes $\frac{1}{2}$ of the GPU memory consumed by LIC-TCM when compressing $2048\times 2048$ images. When compressing $2560\times 2560$ images, our ${}^{++}$ only consumes about $\frac{1}{4}$ of the GPU memory consumed by LIC-TCM. When compressing a $3584\times 3584$ image, the GPU memory consumption of our proposed ${}^{++}$ is still only $12.3$ GB while the LIC-TCM requires $45.95$ GB GPU memory. The curve of the GPU memory consumed by our ${}^{++}$ as the resolution grows is also much flatter.
4.4.2 On Encoding and Decoding Time
@@
4.4.3 On Forward MACs
Compared with ${}^{+}$, our ${}^{++}$, which employs the proposed linear complexity global spatial context modules, has lower MACs. Compared with the recent LIC-TCM (Liu et al., 2023), our ${}^{++}$ demonstrates significantly reduced MACs. Compared with ELIC (He et al., 2022), STF (Zou et al., 2022), and WACNN (Zou et al., 2022), our ${}^{++}$ has higher MACs when the input is high-resolution image. One contributing factor to the higher MACs is the utilization of Cheng’20 (Cheng et al., 2020) as the basis for the transform module in our proposed ${}^{++}$. While the context module in Cheng’20 is simpler than that of ELIC, STF, and WACNN, the overall MACs of Cheng’20 are higher due to its transform modules having higher MACs. Since our ${}^{++}$ employs a modification of analysis transform and synthesis transform of Cheng’20, it results in higher MACs compared to ELIC and WACNN. In addition, since our ${}^{++}$ employed an more advanced entropy model, which leads to higher MACs. However, the entropy model has modest effect on the overall MACs since the input image is down-sampled for four times in analysis transform, which implies that designing more advanced entropy models is more resource-efficient.
4.5 Ablation Studies
4.5.1 Settings
Figure 14: Inter-slice global spatial attention map $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i}_{q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i}_{k}\right)^{\top}$ of Kodim01, Kodim11, Kodim20, Kodim21 from Kodak (Kodak, 1993) dataset, where $i$ is the index of selected slice, $j$ is the index of selected query.
We conduct corresponding ablation studies and evaluate the contribution of each module on Kodak (Kodak, 1993) dataset. Each model is optimized for MSE. We train each model for 2M steps. We use the training strategy in Section 4.1.
4.5.2 Analysis of Channel-wise Context Module
@@
4.5.3 Analysis of Local Spatial Context Module
@@
The vanilla checkerboard context module leads to slight performance degradation (He et al., 2021) compared with pixel-cnn-like serial context modules (Van den Oord et al., 2016; Minnen et al., 2018). The vanilla checkerboard context module contains one convolutional layer which is linear. The other drawback is the fixed kernel weights during inference. In contrast, our proposed checkerboard attention-based local spatial context module is non-linear and incorporates dynamic attention map generation with two-pass decoding, allowing for improved flexibility and adaptability. Our proposed checkerboard attention-based local spatial context module achieves a further reduction of $2.32\%$ in bit-rate compared to vanilla checkerboard context module. Compared with Case 7, which solely incorporates channel-wise context modules, models incorporating both local spatial and channel context modules demonstrate superior performance, further validating the presence of redundancy in the local spatial domain.
4.5.4 Analysis of Intra-Slice Global Context Module
@@
The $i$-th slice is used as an example to illustrate the process. In practice, the computation of $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\hat{\boldsymbol{y}}^{i}_{ac,v}$ in Equation 15 is performed first for linear complexity. However, we can still compute $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i}_{na,q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i}_{ac,k}\right)^{\top}$ as the attention map to validate the ability to capture global dependencies since the $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i}_{na,q}\right)\left[j\right]\times\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i}_{ac,k}\right)^{\top}$ is employed as the implicit similarity metric, where $j$ is the index of selected query. The attention maps of Kodim01, Kodim11, Kodim 20, Kodim21 captured by proposed intra-slice global spatial context module $g_{gc,intra}$ is illustrated in Fig. 13. The checkerboard-like pattern in the attention map arises due to the absence of interactions within the anchor and non-anchor parts. Our model successfully captures distant correlations between the anchor and non-anchor parts, which local context modules are unable to achieve. Although our intra-slice global context module may bear some resemblance to cross-attention models, we focus solely on the interactions within a single slice. We only use the attention map of $\hat{\boldsymbol{y}}^{i-1}$ to predict correlations in $\hat{\boldsymbol{y}}^{i}$. When our proposed global context modules collaborate with local spatial context modules, the overall performance is further improved, underscoring the necessity of global spatial context modules for capturing global correlations and local spatial context modules for capturing local correlations.
4.5.5 Analysis of Inter-Slice Global Context Module